# Improving Outdoor Multi-cell Fingerprinting-based Positioning via Mobile Data Augmentation

### v1

@@ Abstract
Accurate outdoor positioning in cellular networks is hindered by sparse, heterogeneous measurement collections and the high cost of exhaustive site surveys. This paper introduces a lightweight, modular mobile data augmentation framework designed to enhance multi-cell fingerprinting-based positioning using operator-collected minimization of drive test (MDT) records. The proposed approach decouples spatial and radio-feature synthesis: kernel density estimation (KDE) models the empirical spatial distribution to generate geographically coherent synthetic locations, while a k-nearest-neighbor (KNN)-based block produces augmented per-cell radio fingerprints. The architecture is intentionally training-free, interpretable, and suitable for distributed or on-premise operator deployments, supporting privacy-aware workflows. We both validate each augmentation module independently and assess its end-to-end impact on fingerprinting-based positioning using a real-world MDT dataset provided by an Italian mobile network operator across diverse urban and peri-urban scenarios. Results show that the proposed KDE-KNN augmentation consistently improves positioning performance, with the largest benefits in sparsely sampled or structurally complex regions; we also observe region-dependent saturation effects as augmentation increases. The framework offers a practical, low-complexity path to enhance operator positioning services using existing mobile data traces.
@@ I Introduction
Recent years have seen a growing demand for accurate and reliable positioning services in dense urban areas, indoor environments, and under adverse weather conditions, such as overcast skies, where satellite-based systems like the global positioning system (GPS) often suffer from severe multipath propagation, signal blockage, urban canyon effects, and are known to be power-intensive, making it unsuitable for energy-constrained devices commonly used in mobile applications [1].
In these scenarios, multicell fingerprint-based positioning has emerged as a promising approach due to its robustness in non-line-of-sight conditions and the ability to leverage existing cellular infrastructure [2, 3]. The 3rd generation partnership project (3GPP) incorporated fingerprinting-based positioning into LTE networks under Release 12 [4]. This approach estimates a device’s position by comparing its observed radio measurements, such as reference signal received power (RSRP), to a pre-constructed database of location-tagged fingerprints. By leveraging existing cellular infrastructure, fingerprinting provides robust performance in dense urban and indoor environments, where signal blockage and multipath propagation often degrade the accuracy of conventional techniques. However, achieving high positioning accuracy through multi-cell fingerprinting requires large volumes of spatially and radio‐feature‐rich data in order to capture fine‐grained variations across coverage areas; in practice, collecting such dense measurement sets is both costly and time‐consuming for mobile network operators [5], which poses a significant barrier to widespread deployment.
@@
Statistical insufficiency of training data is one of the core limitations of fingerprinting. It manifests in coarse positioning granularity, especially in cell‐dense regions where small geography can exhibit rapid changes in received signal levels, affecting positioning algorithms’ performance. Moreover, collecting exhaustive user traces raises privacy concerns, as these measurements may encode sensitive movement patterns [6]. These challenges motivate the generation of synthetic mobile measurement data that can augment real traces, filling spatial gaps in the fingerprint database while mitigating the need for additional data collection campaigns.
@@
To improve positioning performance in cellular networks, we introduce a mobile data augmentation framework that synthesizes large volumes of realistic radio measurements. Our proposed computationally-light mobile data generator combines kernel density estimation (KDE) for estimating spatial components with non‐parametric regression leveraging k‐nearest‐neighbor (KNN) for multi‐cell RSRP features generation. This methodology aims to produce dense, geographically coherent synthetic mobile data that mimics the statistics of real-world measurements. Although this framework is designed and tested end-to-end for fingerprinting-based positioning, the augmented measurements it generates may also benefit other tasks, including radio-environment map reconstruction, training and validation of learning-based models for coverage/capacity estimation and link-quality prediction, data-efficient handover and beam-management policy learning, and anomaly/fault detection in network monitoring [7].
@@
To train and validate this augmentation engine, we leverage the minimization of drive test (MDT) framework introduced in 3GPP Release 10 [8]. In MDT, user equipments periodically report field measurements, including RSRP, into the network’s operation & maintenance (O&M) system in order to facilitate failure forecasting, troubleshooting, and network optimization [9]. This framework provides a cost-efficient alternative to traditional drive testing and offers a rich source of data for both network management and optimization. Leveraging a real-world, large-scale MDT dataset provided by an Italian MNO, we systematically investigate the positioning performance gains introduced by our synthetic data augmentation pipeline, quantifying improvements in positioning accuracy and robustness across diverse urban scenarios.
I-A Related Work
Data augmentation has emerged as a powerful technique across several domains, including image classification [10, 11] and audio recognition [12], to mitigate data scarcity, enhance generalization, and improve model robustness. In the context of wireless positioning, these challenges are even more pronounced due to the significant cost and manual effort involved in collecting ground-truth fingerprints in real-world environments, motivating the development of augmentation strategies that can synthetically expand datasets while preserving the spatial and statistical characteristics of the underlying radio environment. While established geo-statistical estimators like Kriging explicitly model spatial correlation to produce geo-tagged features at unmeasured locations [13], these approaches often have high sensitivity to model assumptions, rely on the assumption of data stationarity and isotropy, and become computationally intensive for very large datasets.
A recent survey [14] has documented various augmentation strategies developed for WiFi-based indoor fingerprinting, such as spatial interpolation and signal transformations by leveraging generative models. However, these approaches are proven to be effective in dense indoor settings using short-range signal features, while their efficacy in outdoor cellular scenarios, which involve sparse, noisy, and heterogeneous measurements, is not thoroughly investigated yet. Several positioning approaches, which typically rely on signal propagation models [15, 16] or fingerprinting systems supported by specialized hardware deployments [17, 18], can yield accurate results in controlled environments, but they are difficult to scale in large outdoor areas where exhaustive site surveys are impractical and radio propagation is highly variable in space and time. As such, there remains a pressing need for augmentation methods tailored specifically to outdoor cellular positioning challenges. Building on this indoor-focused literature, Gaussian process regression (GPR) has been explored for augmenting WiFi fingerprints in [19] to improve positioning performance. Although initially designed for indoor settings, we nonetheless treat GPR as a meaningful baseline for radio-feature augmentation and include it in our experimental comparisons to quantify its strengths and limitations on outdoor data.
In a recent study, the GUMBLE framework [6] has been proposed as a general-purpose generative model for mobile network data using Bayesian conditional inference. While effective in learning conditional relationships in network data, this framework is not tailored to generate synthetic data to improve positioning accuracy; instead, it focuses on quantifying the aleatoric and epistemic uncertainty of a synthetically generated mobile dataset. Other recent generative frameworks have explored conditioning the signal generation process on auxiliary environmental features such as terrain, cell orientation, and transmission characteristics [20], while other approaches have leveraged satellite imagery or GPS traces to estimate signal quality at unmeasured locations [21]. However, such contextual inputs are rarely available in practice. In contrast with the existing literature, the focus of our work is to develop a data-driven, context-agnostic augmentation framework specifically tailored for wireless positioning in large-scale outdoor settings that requires only geo-tagged dataset without auxiliary environmental metadata. Our approach explicitly targets the generation of synthetic user locations and radio features in a manner that reflects the true characteristics of real-world mobile datasets to enhance positioning accuracy, particularly in urban regions where data scarcity and signal irregularities are most pronounced.
I-B Our Contributions and Paper Organization
@@
  • Unlike previous studies that focused primarily on synthetic data augmentation to improve positioning in controlled or indoor scenarios, this work addresses the augmentation of real MDT data provided by an Italian MNO in four large-scale outdoor settings, which introduces practical challenges such as sparse spatial sampling, heterogeneous and partial physical cell identity (PCI) visibility, measurement noise, and strong temporal/spatial variability due to user mobility and wide-area propagation effects.
@@
  • We empirically validate each module of the augmentation pipeline in isolation. For the spatial augmentation stage, we compare KDE against parametric and learned alternatives ( Gaussian mixture model (GMM), generative adversarial network (GAN), and normalizing flow (NF)) using a multivariate two-sample Kolmogorov-Smirnov (KS) test to quantify distributional agreement with real MDT locations. For the radio augmentation stage, we benchmark the proposed KNN approach against random forest (RF) and GPR with both squared exponential (SE) and rational quadratic (RQ) kernels, assessing prediction quality in terms of mean absolute error (MAE) on per-PCI RSRP values.
  • We assess the end-to-end impact of augmentation on multi-cell fingerprinting-based positioning by comparing the proposed KDE-KNN pipeline against representative two-stage baselines (KDE-RF, KDE-GPR (SE), KDE-GPR (RQ)) by means of average positioning errors over multiple tests.
@@
The remainder of the paper is structured as follows. Sec. II introduces the proposed mobile data augmentation pipeline, describing the modular components for the spatial and radio features generation. Sec. III discusses both the multi-cell fingerprinting-based positioning problem and introduces the four real-world MDT-based reference scenarios Sec. IV presents the numerical results in terms of (i) independent performance of each stage of the modular framework, (ii) performance benefits of augmentation to fingerprinting-based positioning in the considered scenarios, and (iii) analysis on performance scaling with respect to the amount of synthetic data generated. Finally, Sec. V concludes the paper.
II Mobile Data Augmentation
Starting from a set of geo-tagged MDT records collected by the MNO , the proposed pipeline proceeds in two stages. First, KDE is used to estimate the empirical spatial distribution of measurements and to draw synthetic sample locations. Second, for each synthetic location, a multi-cell RSRP vector is predicted via a KNN operating on the original MDT fingerprints. This produces augmented MDT data that can be integrated into fingerprinting-based positioning algorithms. The overall flow is illustrated in Fig. 1. MDT data structure is discussed in Sec. II-A, while the two-stage architecture details are given in Sec. II-B and Sec. II-C.
II-A MDT Data
$\text{RSRP}=\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{14}P_{\text{RE},ik},$ (1)
@@ II-A MDT Data
II-B Spatial Features Augmentation via Density Estimation
@@
Specifically, we adopt KDE, a non-parametric technique, motivated by its simplicity, interpretability, and suitability for modeling the spatial distribution of users. In Sec. IV-A, we further validate this choice by comparing performance using the multivariate KS test against parametric baselines such as GANs, NFs, and GMMs. Specifically, KDE estimates the underlying probability density function (PDF) by placing a kernel $K(\mathbf{p}-\mathbf{p}_{i},h)$ centered at each training point $\mathbf{p}_{i}\in\mathcal{D}$, with $h$ representing the kernel bandwidth. This yields an interpretable PDF estimate $f(\mathbf{p},\mathcal{D},h):\mathbb{R}^{2}\rightarrow\mathbb{R}^{+}$, constructed directly from the empirical data, capturing multi-modal structures inherent in the distribution [31]. Formally, the KDE is defined as
@@
We select the Gaussian kernel based on domain-specific considerations. MDT location data are typically obtained via GPS at the user side, and the associated measurement noise is typically modeled as Gaussian. Therefore, using a Gaussian kernel with a tunable bandwidth aligns naturally with the assumed uncertainty in user positioning. Intuitively, KDE assumes that regions with higher sample density are more likely to generate future observations. Since KDE is an unsupervised method, the bandwidth $h$ is treated as a hyperparameter and tuned via ERM. Specifically, we minimize the negative log-likelihood on a validation subset $\mathcal{D}^{\prime}=\{\mathbf{p}_{i}^{\prime}\}_{i=1}^{m^{\prime}}$, that is
@@
II-C Radio Features Augmentation via Non-Parametric Regression
Radio features augmentation can be addressed with a range of approaches; representative examples include neural networks [6], RFs [32], and GPRs [19], which have been used for related tasks in literature. In this work, we adopt KNN, which is a simple, interpretable non-parametric procedure that transfers measured per-PCI radio features from nearby reference locations. In Sec. IV-B, we further validate this choice by comparing performance using MAE on radio feature prediction against baselines such as RFs and GPRs.

### v2

@@ Abstract
Accurate outdoor positioning in cellular networks is hindered by sparse, heterogeneous measurement collections and the high cost of exhaustive site surveys. This paper introduces a lightweight, modular mobile data augmentation framework designed to enhance multi-cell fingerprinting-based positioning using operator-collected minimization of drive test (MDT) records. The proposed approach decouples spatial and radio-feature synthesis: kernel density estimation (KDE) models the empirical spatial distribution to generate geographically coherent synthetic locations, while a k-nearest-neighbor (KNN)-based block produces augmented per-cell radio fingerprints. The architecture is intentionally training-free, interpretable, and suitable for distributed or on-premise operator deployments, supporting privacy-aware workflows. We both validate each augmentation module independently and assess its end-to-end impact on fingerprinting-based positioning using a real-world MDT dataset provided by an Italian mobile network operator across diverse urban and peri-urban scenarios. Results show that the proposed KDE-KNN augmentation consistently improves positioning performance with respect to state-of-the-art approaches, reducing the median positioning error by up to 30% in the most sparsely sampled or structurally complex regions. We also observe region-dependent saturation effects, which emerge most rapidly in scenarios with high user density where the information gain from additional synthetic samples quickly diminishes. Overall, the framework offers a practical, low-complexity path to enhance operator positioning services using existing mobile data traces.
@@ I Introduction
Recent years have seen a growing demand for accurate and reliable positioning services in dense urban areas and indoor environments [47]. While satellite-based systems like the global positioning system (GPS) remain the gold standard for global positioning, their utility in operational mobile networks is frequently constrained by practical factors, as they remain power-intensive components [19]. Consequently, mobile operating systems and user policies often disable or duty-cycle satellite tracking to conserve battery life. In these scenarios, multicell fingerprint-based positioning has emerged as a promising approach due to its robustness in non-line-of-sight conditions. However, statistical insufficiency of training data is one of the core limitations of fingerprinting. It manifests in coarse positioning granularity, especially in cell‐dense regions where small geography can exhibit rapid changes in received signal levels, affecting positioning algorithms’ performance. Moreover, collecting exhaustive user traces raises privacy concerns, as these measurements may encode sensitive movement patterns [37]. These challenges motivate the generation of synthetic mobile measurement data that can augment real traces, filling spatial gaps in the fingerprint database while mitigating the need for additional data collection campaigns.
@@
To improve positioning performance in cellular networks, we introduce a mobile data augmentation framework that synthesizes large volumes of realistic radio measurements. Our proposed computationally-light mobile data generator combines kernel density estimation (KDE) for estimating spatial components with non‐parametric regression leveraging k‐nearest‐neighbor (KNN) for multi‐cell reference signal received power (RSRP) features generation. This methodology aims to produce dense, geographically coherent synthetic mobile data that mimics the statistics of real-world measurements. Although this framework is designed and tested end-to-end for fingerprinting-based positioning, the augmented measurements it generates may also benefit other tasks, including radio-environment map reconstruction, training and validation of learning-based models for coverage/capacity estimation and link-quality prediction, data-efficient handover and beam-management policy learning, and anomaly/fault detection in network monitoring [5].
@@
In this work, by leveraging real-world, large-scale minimization of drive test (MDT) data provided by an Italian mobile network operator (MNO), we systematically investigate the positioning performance gains introduced by our proposed synthetic data augmentation pipeline, quantifying improvements in positioning accuracy and robustness across diverse urban scenarios.
I-A Contributions and Paper Organization
@@
  • Unlike previous studies that focused primarily on synthetic data augmentation to improve positioning in controlled or indoor scenarios, this work addresses the practical challenges inherent to operational outdoor environments. We validate the proposed framework on real MDT data provided by an Italian MNO in four large-scale settings, demonstrating its capability to handle sparse spatial sampling, heterogeneous and partial physical cell identity (PCI) visibility, measurement noise, and strong temporal/spatial variability typical of user mobility and wide-area propagation effects.
@@
  • We empirically validate each module of the augmentation pipeline in isolation. For the spatial augmentation stage, we compare KDE against parametric and learned alternatives ( Gaussian mixture model (GMM), generative adversarial network (GAN), and normalizing flow (NF)) using a multivariate two-sample Kolmogorov-Smirnov (KS) test to quantify distributional agreement with real MDT locations. For the radio augmentation stage, we benchmark the proposed KNN approach against random forest (RF), Gaussian process regression (GPR) with both squared exponential (SE) and rational quadratic (RQ) kernels, multilayer perceptron (MLP), conditional generative adversarial network (CGAN), rectified flow (ReFlow) (diffusion model), and NF. We assess prediction quality in terms of median absolute error (MedAE) on per-PCI RSRP values.
  • We assess the end-to-end impact of augmentation on multi-cell fingerprinting-based positioning by comparing the proposed KDE-KNN pipeline against representative two-stage baselines (KDE-RF, KDE-GPR) and advanced deep learning architectures (KDE-MLP, KDE-CGAN, KDE-ReFlow, and KDE-NF) by means of average positioning errors over multiple tests.
@@
  • To facilitate reproducibility and support further research in the field, we make the source code of the proposed framework publicly available. The repository can be accessed at: https://github.com/wilabcnit/Mobile-Data-Augmentation-Positioning.
@@
The remainder of the paper is structured as follows. Sec. II provides the background on fingerprinting and mobile data collection. Sec. III reviews the related work. Sec. IV introduces the proposed mobile data augmentation pipeline, describing the modular components for the spatial and radio features generation. Sec. V discusses both the multi-cell fingerprinting-based positioning problem and introduces the four real-world MDT-based reference scenarios Sec. VI presents the numerical results in terms of (i) independent performance of each stage of the modular framework, (ii) performance benefits of augmentation to fingerprinting-based positioning in the considered scenarios, and (iii) analysis on performance scaling with respect to the amount of synthetic data generated. Finally, Sec. VII concludes the paper.
II Background
II-A Fingerprinting in Cellular Networks
Fingerprint-based positioning leverages existing cellular infrastructure and has become a standard for robust positioning in dense urban and indoor environments [16, 13]. The 3rd generation partnership project (3GPP) incorporated fingerprinting-based positioning into LTE networks under Release 12 [1]. This approach estimates a device’s position by comparing its observed radio measurements, such as RSRP, to a pre-constructed database of location-tagged fingerprints. By leveraging existing cellular infrastructure, fingerprinting provides robust performance where signal blockage and multipath propagation often degrade the accuracy of conventional techniques. However, achieving high positioning accuracy through multi-cell fingerprinting requires large volumes of spatially and radio‐feature‐rich data in order to capture fine‐grained variations across coverage areas; in practice, collecting such dense measurement sets is both costly and time‐consuming for MNOs [7], which poses a significant barrier to widespread deployment.
II-B Minimization of Drive Tests (MDT)
To enable data-driven optimization, we leverage the MDT framework introduced in 3GPP Release 10 [4]. In MDT, user equipments periodically report field measurements, including RSRP, into the network’s operation & maintenance (O&M) system in order to facilitate failure forecasting, troubleshooting, and network optimization [36, 30]. While this framework provides a rich, cost-efficient alternative to traditional drive testing, it is subject to the operational constraints of commercial user devices. Notably, in the commercial MDT dataset analyzed in this work, valid GPS coordinates are present in fewer than 4% of the reported samples, as devices frequently disable satellite positioning to conserve power. This sparsity necessitates the development of augmentation strategies that can effectively utilize the limited available geolocated data.
III Related Work
Data augmentation has emerged as a powerful technique across several domains, including image classification [20, 24] and audio recognition [35], to mitigate data scarcity, enhance generalization, and improve model robustness. In the context of wireless positioning, these challenges are even more pronounced due to the high cost and manual effort involved in collecting ground-truth fingerprints in real-world environments, motivating the development of augmentation strategies that can synthetically expand datasets while preserving the spatial and statistical characteristics of the underlying radio environment. While established geo-statistical estimators like Kriging explicitly model spatial correlation to produce geo-tagged features at unmeasured locations [28], these approaches often have high sensitivity to model assumptions, rely on the assumption of data stationarity and isotropy, and become computationally intensive for very large datasets.
A recent survey [11] has documented various augmentation strategies developed for WiFi-based indoor fingerprinting, such as spatial interpolation and signal transformations by leveraging generative models. However, these approaches are proven to be effective in dense indoor settings using short-range signal features, while their efficacy in outdoor cellular scenarios, which involve sparse, noisy, and heterogeneous measurements, is not thoroughly investigated yet. Transferring these methods to the outdoor domain is non-trivial due to distinct operational characteristics: unlike static indoor layouts, wide-area cellular networks exhibit significant variability in macro-cell density and coverage geometry, alongside practical anomalies such as incomplete PCI visibility and unpredictable user mobility patterns [23].
Several positioning approaches, which typically rely on signal propagation models [26, 46] or fingerprinting systems supported by specialized hardware deployments [18, 27], can yield accurate results in controlled environments, but they are difficult to scale in large outdoor areas where exhaustive site surveys are impractical and radio propagation is highly variable in space and time. Furthermore, operational outdoor networks are subject to continuous temporal drift driven by environmental non-stationarity (e.g., seasonal changes), traffic-dependent interference, and long-term network evolution, all of which degrade the longevity and reliability of static fingerprint databases [12]. As such, there remains a pressing need for augmentation methods tailored specifically to outdoor cellular positioning challenges. Building on this indoor-focused literature, GPR has been explored for augmenting WiFi fingerprints in [39] to improve positioning performance. Although initially designed for indoor settings, we nonetheless treat GPR as a meaningful baseline for radio-feature augmentation and include it in our experimental comparisons to quantify its strengths and limitations on outdoor data.
It is worth noting that radio features can also be predicted using deterministic approaches, such as ray tracing (RT) [15, 41], or empirical propagation models [48]. While RT can provide accurate channel estimation, it requires precise 3D environmental models (digital twins) and detailed infrastructure metadata (e.g., antenna patterns, material properties), which are computationally expensive to process and rarely available for large-scale operational networks. Similarly, standard empirical models often lack the granularity to capture local shadowing effects essential for fingerprinting. Furthermore, both approaches typically model the physical channel rather than the reported measurement data, without accounting for device heterogeneity, body-loss, and measurement noise present in real MDT traces. In contrast, data-driven approaches allow for feature augmentation that implicitly learns these complex statistical distributions directly from the data.
In a recent study, the GUMBLE framework [37] has been proposed as a general-purpose generative model for mobile network data using Bayesian conditional inference. While effective in learning conditional relationships in network data, this framework is not tailored to generate synthetic data to improve positioning accuracy; instead, it focuses on quantifying the aleatoric and epistemic uncertainty of a synthetically generated mobile dataset. Other recent generative frameworks have explored conditioning the signal generation process on auxiliary environmental features such as terrain, cell orientation, and transmission characteristics [38], while other approaches have leveraged satellite imagery or GPS traces to estimate signal quality at unmeasured locations [40]. However, such contextual inputs are rarely available in practice. Recent literature supports the view that complex generative models may not be optimal for all scenarios. For instance, the survey in [31] indicates that for low-dimensional, non-smooth radio data where network geometry is unknown, interpolation-based approaches can offer superior robustness compared to deep learning methods. Building on this insight, and in contrast with the existing literature, the focus of our work is to develop a data-driven, context-agnostic augmentation framework specifically tailored for positioning in large-scale outdoor settings. Our proposed method requires only geo-tagged datasets without auxiliary environmental metadata, and enables the generation of synthetic user locations and radio features in a manner that reflects the true characteristics of real-world mobile datasets to enhance positioning accuracy, particularly in urban regions where data scarcity and signal irregularities are most pronounced.
@@ III Related Work
IV Mobile Data Augmentation
Starting from a set of geo-tagged MDT records collected by the MNO, the proposed pipeline proceeds in two stages. First, KDE is used to estimate the empirical spatial distribution of measurements and to draw synthetic sample locations. Second, for each synthetic location, a multi-cell RSRP vector is predicted via a KNN operating on the original MDT fingerprints. This produces augmented MDT data that can be integrated into fingerprinting-based positioning algorithms. The overall flow is illustrated in Fig. 1. MDT data structure is discussed in Sec. IV-A, while the two-stage architecture details are given in Sec. IV-B and Sec. IV-C.
IV-A MDT Data
$\text{RSRP}=\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{14}P_{\text{RE},ik},$ (1)
@@ IV-A MDT Data
IV-B Spatial Features Augmentation via Density Estimation
@@
Specifically, we adopt KDE, a non-parametric technique, motivated by its simplicity, interpretability, and suitability for modeling the spatial distribution of users. Unlike black-box generative models, KDE offers a transparent decision mechanism where the estimated density is directly traceable to the proximity of specific training examples. In Sec. VI-A, we further validate this choice by comparing performance using the multivariate KS test against parametric baselines such as GANs, NFs, and GMMs. Specifically, KDE estimates the underlying probability density function (PDF) by placing a kernel $K(\mathbf{p}-\mathbf{p}_{i},h)$ centered at each training point $\mathbf{p}_{i}\in\mathcal{D}$, with $h$ representing the kernel bandwidth. This yields an interpretable PDF estimate $f(\mathbf{p},\mathcal{D},h):\mathbb{R}^{2}\rightarrow\mathbb{R}^{+}$, constructed directly from the empirical data, capturing multi-modal structures inherent in the distribution [44]. Formally, the KDE is defined as
@@
We select the Gaussian kernel based on domain-specific considerations. MDT location data are typically obtained via GPS at the user side, and the associated measurement noise is typically modeled as Gaussian. Therefore, using a Gaussian kernel with a tunable bandwidth aligns naturally with the assumed uncertainty in user positioning. This reinforces the model’s interpretability: the predicted density at any point $\mathbf{p}$ is not an abstract score, but a cumulative measure of how many training users were physically located within the noise radius $h$ of that point. Intuitively, KDE assumes that regions with higher sample density are more likely to generate future observations. Since KDE is an unsupervised method, the bandwidth $h$ is treated as a hyperparameter and tuned via ERM. Specifically, we minimize the negative log-likelihood on a validation subset $\mathcal{D}^{\prime}=\{\mathbf{p}_{i}^{\prime}\}_{i=1}^{m^{\prime}}$, that is
@@
IV-C Radio Features Augmentation via Non-Parametric Regression
Radio features augmentation can be addressed with a range of approaches; representative examples include neural networks [37], RFs [42], and GPRs [39], which have been used for related tasks in literature. In this work, we adopt KNN, which is a simple, interpretable non-parametric procedure that transfers measured per-PCI radio features from nearby reference locations. In Sec. VI-B, we further validate this choice by comparing performance using MedAE on radio feature prediction against machine learning approaches (RF and GPR) and deep learning approaches (MLP, CGAN, ReFlow, and NF).